Time table for WS 07/08

08:00 - 09:00 09:00 - 10:00 10:00 - 11:00 11:00 - 12:00 12:00 - 13:00 13:00 - 14:00 14:00 - 15:00 15:00 - 16:00 16:00 - 17:00 Monday Tuesday Wednesday Thursday

Friday

Image Processing and Computer Vision

Weickert

E13 002

Image Processing and Computer Vision (Tutorial)

Zimmer E13 016

Introduction to Image Acquisition Methods

Weickert E13 001

Image Processing and Computer Vision

Weickert E13 002